OpenStudy (anonymous):

i need help!!! wat is equivilant to i to the power of 903

OpenStudy (hoblos):

i^903 = (i^2)^451 * i = (-1)^451 *i = -i

OpenStudy (anonymous):

thankyou soo much

OpenStudy (hoblos):

my pleasure XD